Suppose a firm faces the production function Q=4L+12L^2-2L^3



find the maximum points of AP and MP curves

1
Expert's answer
2022-06-06T08:44:33-0400

"Q=4L+12L^2-2L^3,"

"Q'=4+24L-6L^2,"

"Q'=0,"

"L_1=-0.16,"

"L_2=4.16,"

"MP=Q(L_1)=-0.32,"

"AP=Q(L_2)=80.32."